OpenStudy (anonymous):

the speed of a passenger train is 16mph faster than the speed of a freight train. The passenger train travels 290 miles in the same time it takes the freight train to travel 210 miles. Find the speed of each train

OpenStudy (anonymous):

freight speed = x passenger speed = x+16 x/210 = (x+16)/290 290x=(210)(x+16) 290x=210x+3360 80x=3360 x=42 speed of freight is 42mph speed of passenger is 42+16 or 58 mph. Proof: 210/42 = 5 and 290/(42+16)=5 both are equal so this proves answer of 42 mph and 58 mph.